OpenStudy (anonymous):

How to solve sin(3x)= 1/2 ?

OpenStudy (nubeer):

replace 3x=y sin(Y) = 0.5 now find the angels for y.. can u do that?

OpenStudy (anonymous):

i got x = pi/18

OpenStudy (anonymous):

is that the right answer?

OpenStudy (nubeer):

yes .. but this is one angle.. you have to find the other angle too.

OpenStudy (anonymous):

How do I find that? The problem doesn't specify a restriction, so do I have to find the infinite number of angles?

OpenStudy (nubeer):

no.. not infinite.. it's like this.. sin(3x) =0.5 you have to look after the equal sign whether it's positive or negative , in this case it's positive.. so you have to see in which quadrants the value of sin is positive.. 1st and 2nd quadrants.. in both of these the value of sin is positive..

OpenStudy (nubeer):

3x = pi/6 so pi/6 is in first quadrant.. so you have to find the angel for second quadrant too.. that you can find by pi-pi/6 3x = pi , 5pi/6 now solve it

OpenStudy (anonymous):

Don't you mean 3x = pi/6, 5pi/6?

OpenStudy (anonymous):

I got pi/18 and 5pi/18

OpenStudy (nubeer):

yup.. exactly.

OpenStudy (nubeer):

i was just explaining incase you didn't knew...

OpenStudy (anonymous):

Okay, thank you! I have one more question, could you help me with that?

OpenStudy (nubeer):

yup sure..

OpenStudy (anonymous):

tan2x = 1 in the interval [-π/2, 3π/2]

OpenStudy (nubeer):

ok.. so have you started solving this?

OpenStudy (anonymous):

I started with 2x = 1, so 2x = π/4 and 5π/4. I just don't get what the restriction means

OpenStudy (nubeer):

the restriction tells that you have to tell all the angles between that range.. for that you have to know in which quadrants tan is positive.

OpenStudy (anonymous):

Quadrants I and III

OpenStudy (nubeer):

yes and any of these quadrants comes between the range defined in question?

OpenStudy (anonymous):

What does [-π/2, 3π/2] mean? It makes one revolution, but doesn't that mean I can't say π/4 and 5π/4?

OpenStudy (nubeer):



OpenStudy (nubeer):